# Heterogeneous Multi-agent Zero-Shot Coordination by Coevolution

### v2

@@ Abstract
Generating agents that can achieve zero-shot coordination (ZSC) with unseen partners is a new challenge in cooperative multi-agent reinforcement learning (MARL). Recently, some studies have made progress in ZSC by exposing the agents to diverse partners during the training process. They usually involve self-play when training the partners, implicitly assuming that the tasks are homogeneous. However, many real-world tasks are heterogeneous, and hence previous methods may be inefficient. In this paper, we study the heterogeneous ZSC problem for the first time and propose a general method based on coevolution, which coevolves two populations of agents and partners through three sub-processes: pairing, updating and selection. Experimental results on various heterogeneous tasks highlight the necessity of considering the heterogeneous setting and demonstrate that our proposed method is a promising solution for heterogeneous ZSC tasks.
@@ I Introduction
In many real-world tasks, AI and humans are, however, usually heterogeneous [17, 47], as shown in Figure 1. That is, agents and partners have different advantages (i.e., the efficiency of doing something) or skills (i.e., the ability to do something). For example, the robotic arm can lift heavy objects that human workers cannot, and workers can perform more flexible operations than the robotic arm. The heterogeneous setting has greater applicability and can recover homogeneous policies as a result if necessary [18]. Though heterogeneous tasks are widely available and important, it is nontrivial to train heterogeneous agents, especially in the ZSC setting. Previous self-play methods mainly focus on homogeneous settings, which may not capture the cooperative behaviors between agents and human partners well, thus being inefficient for heterogeneous ZSC tasks. Thus, a natural question is raised: Can we generate agents without human data that can coordinate with novel and heterogeneous humans efficiently?
@@ III-B Diversity in RL
A series of works have explored the benefit of diversity in RL. For example, diversity has been found helpful in fast adaptation [8], RL with sparse or deceptive rewards [7], open-ended learning [48, 41], unsupervised environment designing [32] and ZSC [26, 42, 54]. There are various definitions of diversity [19, 34, 26], and they are summarized in Liu et al. [22].
@@ IV-B Updating
where $\mathcal{A}_{\pi}$ denotes the common action space of the policies, and $\bar{\pi}(\cdot\mid s_{t})=\frac{1}{n}\sum^{n}_{i=1}\pi^{(i)}(\cdot\mid s_{t})$. Then, for each policy $\pi$ representing an agent or partner, the objective function is defined as
@@ V Experiments
In this section, we first introduce the environments used for our experiments and the experimental settings. Then, we answer the following four research questions (RQs): 1) Is the heterogeneity worth considering? 2) How does MAZE perform compared with recently proposed algorithms? 3) How do different components of MAZE influence its performance? 4) Can MAZE coordinate well with real human partners? Finally, we provide some additional results, including the influence of larger network, visualization of partner diversity, running time analysis of MAZE, and comparison between homogeneous and heterogeneous environments.
@@ V-B Experimental settings
SP and PP
We employ the commonly used implementation of SP and PP [^1] https://github.com/HumanCompatibleAI/human_aware_rl/tree/neurips2019.
@@
The implementation refers to the official demo. [^2] https://colab.research.google.com/drive/1yIiw0Qs_d3O_JzGxtyvxB--jJ2M0h3Ki The diversity coefficient $\alpha$ is set to $4$.
@@
We maintain $5$ self-play partners initialized with $5$ different seeds (i.e., 1000–5000), respectively. Besides, we save the first checkpoint (i.e., a randomly initialized partner), the middle checkpoint (i.e., a half-trained partner), and the last checkpoint (i.e., a fully trained partner) to obtain the diverse population of partners.
@@
We use the reference implementation of the authors for MEP [^3] https://github.com/ruizhaogit/maximum_entropy_population_based_training and employ the default parameter settings (i.e., weight of the population entropy reward is $0.04$ for the FC layout and $0.01$ for other layouts).
@@ V-C RQ1: Is the heterogeneity worth considering?
We first investigate if considering heterogeneity is valuable during the training process. To examine the importance of considering heterogeneity more purely, we implement the simplest variant of MAZE, i.e., V-MAZE, which uses the fixed pairing strategy, updates the policies without considering the diversity term in the objective function, and does not use the archive for partner selection. That is, the only difference between V-MAZE and SP is that V-MAZE maintains two separate populations of agents and partners. We compare V-MAZE with SP [39] and PP [16], which use self-play and single-population play, respectively.
@@
Table II: The reward (mean±$\pm$±std.) achieved by the compared algorithms when testing with different partners on different layouts. For each combination of layout and partner, the largest reward is bolded. The symbols ‘$+$’, ‘--’ and $\approx$’ indicate that the result is significantly superior to, inferior to, and almost equivalent to MAZE, respectively, according to the Wilcoxon rank-sum test with significance level 0.05.
Layout Partner SP PP TrajeDi FCP MEP MAZE
H-CR Random 3.8 $\pm$ 2.1 -- 5.0 $\pm$ 1.5 -- 5.5 $\pm$ 0.7 -- 6.4 $\pm$ 1.1 $\approx$ 6.5 $\pm$ 0.7 $\approx$ 7.0 $\pm$ 0.7
Self-Play 40.4 $\pm$ 9.5 -- 36.8 $\pm$ 6.4 -- 87.6 $\pm$ 9.0 $\approx$ 121.6 $\pm$ 12.9 $+$ 85.8 $\pm$ 7.5 $\approx$ 97.5 $\pm$ 14.6
MAZE 100.0 $\pm$ 14.4 -- 143.7 $\pm$ 17.1 -- 153.8 $\pm$ 16.5 -- 140.4 $\pm$ 15.5 -- 129.8 $\pm$ 14.9 -- 183.7 $\pm$ 11.9
Human Proxy 61.4 $\pm$ 10.3 -- 73.6 $\pm$ 10.3 -- 81.0 $\pm$ 13.3 -- 101.2 $\pm$ 15.5 $\approx$ 90.2 $\pm$ 10.0 -- 118.6 $\pm$ 13.2
AA Random 30.6 $\pm$ 7.6 -- 36.0 $\pm$ 9.1 -- 49.3 $\pm$ 11.9 -- 50.4 $\pm$ 7.4 -- 63.5 $\pm$ 12.5 $\approx$ 74.4 $\pm$ 15.5
Self-Play 172.6 $\pm$ 10.4 $+$ 204.8 $\pm$ 10.8 $+$ 180.5 $\pm$ 15.5 $+$ 171.0 $\pm$ 11.5 $+$ 190.3 $\pm$ 10.1 $+$ 142.6 $\pm$ 20.3
MAZE 213.8 $\pm$ 17.0 -- 223.0 $\pm$ 10.1 -- 247.0 $\pm$ 27.1 -- 236.0 $\pm$ 12.9 -- 120.0 $\pm$ 12.4 -- 315.5 $\pm$ 10.1
Human Proxy 38.5 $\pm$ 10.0 -- 43.9 $\pm$ 2.9 -- 50.2 $\pm$ 6.0 -- 38.4 $\pm$ 4.2 -- 92.0 $\pm$ 29.5 $\approx$ 112.0 $\pm$ 19.1
AA-2 Random 28.0 $\pm$ 6.7 -- 29.8 $\pm$ 7.0 -- 45.5 $\pm$ 10.1 $\approx$ 43.5 $\pm$ 10.5 $\approx$ 58.4 $\pm$ 14.8 $\approx$ 56.5 $\pm$ 9.3
Self-Play 52.4 $\pm$ 10.4 -- 79.6 $\pm$ 19.4 -- 100.8 $\pm$ 12.5 -- 121.2 $\pm$ 15.5 $\approx$ 108.0 $\pm$ 29.9 $\approx$ 130.6 $\pm$ 18.5
MAZE 132.6 $\pm$ 29.7 -- 152.8 $\pm$ 24.6 -- 170.5 $\pm$ 11.1 -- 156.0 $\pm$ 15.3 -- 124.6 $\pm$ 29.3 -- 381.4 $\pm$ 13.7
Human Proxy 39.4 $\pm$ 6.2 -- 33.8 $\pm$ 9.3 -- 48.0 $\pm$ 6.9 -- 55.8 $\pm$ 15.2 -- 79.0 $\pm$ 11.2 -- 111.5 $\pm$ 13.7
FC Random 3.3 $\pm$ 2.2 -- 3.0 $\pm$ 0.6 -- 6.0 $\pm$ 0.7 -- 5.1 $\pm$ 0.7 -- 6.0 $\pm$ 1.0 -- 7.3 $\pm$ 0.4
Self-Play 74.6 $\pm$ 17.6 -- 76.8 $\pm$ 25.2 -- 80.8 $\pm$ 22.4 -- 79.5 $\pm$ 22.4 -- 82.3 $\pm$ 22.6 -- 147.0 $\pm$ 27.1
MAZE 96.5 $\pm$ 16.0 -- 92.6 $\pm$ 11.6 -- 115.6 $\pm$ 25.7 -- 119.8 $\pm$ 10.9 -- 95.2 $\pm$ 15.7 -- 164.0 $\pm$ 27.0
Human Proxy 21.3 $\pm$ 2.4 $\approx$ 28.1 $\pm$ 13.3 $\approx$ 25.4 $\pm$ 6.8 $\approx$ 26.9 $\pm$ 7.4 $\approx$ 28.6 $\pm$ 5.3 $\approx$ 25.0 $\pm$ 5.1
+//+/-$\approx$+ / - / ≈ 1/14/1 1/14/1 1/11/3 2/10/4 1/8/7
Average Rank 5.38 4.50 3.19 3.25 3.00 1.69
@@
The training curves are shown in Figure 6, reflecting the change in the average reward of the agents during the training phase. In all the six heterogeneous environments, V-MAZE achieves better performance clearly, showing the necessity of considering the heterogeneity and distinguishing the two players explicitly. Besides, we use a larger scale network to train SP and PP agents on the heterogeneous layout AA in Section V-G. With more generations (i.e., almost three times of V-MAZE), SP and PP still perform worse than V-MAZE. This indicates that the heterogeneous skills of both players cannot be mastered well at the same time even with enhanced representation ability and more data, further strengthening our conclusions.

### v3

@@ Abstract
Generating agents that can achieve zero-shot coordination (ZSC) with unseen partners is a new challenge in cooperative multi-agent reinforcement learning (MARL). Recently, some studies have made progress in ZSC by exposing the agents to diverse partners during the training process. They usually involve self-play when training the partners, implicitly assuming that the tasks are homogeneous. However, many real-world tasks are heterogeneous, and hence previous methods may be inefficient. In this paper, we study the heterogeneous ZSC problem for the first time and propose a general method based on coevolution, which coevolves two populations of agents and partners through three sub-processes: pairing, updating and selection. Experimental results on various heterogeneous tasks highlight the necessity of considering the heterogeneous setting and demonstrate that our proposed method is a promising solution for heterogeneous ZSC tasks. To the best of our knowledge, we are the first to underscore the significance of the heterogeneous ZSC tasks and to introduce an effective framework for addressing it.
@@ I Introduction
In many real-world tasks, AI and humans are, however, usually heterogeneous [20, 53], as shown in Figure 1. That is, agents and partners have different advantages (i.e., the efficiency of doing something) or skills (i.e., the ability to do something). For example, in factory production tasks, AI agents (e.g., robotic arms) can lift heavy objects, while human workers can perform more flexible operations than the robotic arm. In disaster search and rescue missions, AI agents (e.g., drones and quadruped robots) exhibit faster mobility and are better suited to explore dangerous, hard-to-reach areas, while human doctors have slower speed but can better treat the injured. The heterogeneous setting has greater applicability and can recover homogeneous policies as a result if necessary [21]. Though heterogeneous tasks are widely available and important, it is nontrivial to train heterogeneous agents, especially in the ZSC setting. Previous self-play methods mainly focus on homogeneous settings, which may not capture the cooperative behaviors between agents and human partners well, thus being inefficient for heterogeneous ZSC tasks. There is a significant gap between the demands of real-world applications and the deficiencies in heterogeneous algorithm research. Thus, a natural question is raised: Can we generate agents without human data that can coordinate with novel and heterogeneous humans efficiently?
@@
Our contributions are three-folds:
  • Introduction of a Novel Problem Setting: We pioneer the exploration of heterogeneous ZSC. This paper is the first to systematically identify the complexities and necessities of addressing heterogeneity in ZSC, demonstrating its prevalence and significance in many real-world scenarios.
  • Development of an Effective Framework: We introduce MAZE, a novel framework specifically designed for heterogeneous ZSC challenges. Unlike traditional approaches that rely on homogeneous self-play, MAZE innovatively employs a coevolutionary strategy involving two distinct populations of agents and partners. This method not only enhances the adaptability and robustness of AI agents when coordinating with novel, diverse human partners but also effectively addresses the inherent challenges posed by the heterogeneity of capabilities and roles.
  • Comprehensive Empirical Validation and Analysis: Through rigorous experimentation in the Overcooked environment and a custom-designed grid-world setting, FillInTheGrid, we provide a thorough empirical evaluation of MAZE. Our analysis demonstrates the superiority of MAZE over existing methods in the heterogeneous ZSC tasks.
@@ III-B Diversity in RL
A series of works have explored the benefit of diversity in RL. For example, diversity has been found helpful in fast adaptation [8, 42], robust training [59, 61], sparse or deceptive reward [7, 22], open-ended learning [54, 47], unsupervised environment designing [36, 12] and ZSC [31, 48, 62]. There are various definitions of diversity [23, 38, 31], and they are summarized in Liu et al. [27].
@@ IV-B Updating
where $\mathcal{A}_{\pi}$ denotes the common action space of the policies, and $\bar{\pi}(\cdot\mid s_{t})=\frac{1}{n}\sum^{n}_{i=1}\pi^{(i)}(\cdot\mid s_{t})$. Based on the JSD measure Eq. (1), we use policy $\pi$ to represent an agent or a partner, whose objective function can be defined as
@@ V Experiments
In this section, we first introduce the environments used for our experiments and the experimental settings. Then, we answer the following four research questions (RQs): 1) Is the heterogeneity worth considering? 2) How does MAZE perform compared with recently proposed algorithms? 3) How do different components of MAZE influence its performance? 4) Can MAZE coordinate well with real human partners? Finally, we provide some additional results, including the influence of larger network, visualization of partner diversity, running time analysis of MAZE, and comparison between homogeneous and heterogeneous environments. Our code is available at https://github.com/lamda-bbo/maze.
@@ V-B Experimental settings
SP
is a baseline in ZSC, which makes the agent pair with a copy of itself to interact with the environment and update. We employ the commonly used implementation of SP [^1] https://github.com/HumanCompatibleAI/human_aware_rl/tree/neurips2019.
PP
is another baseline in ZSC, which maintains a population of agents that pair with each other. We employ the commonly used implementation of PP [^2] https://github.com/HumanCompatibleAI/human_aware_rl/tree/neurips2019.
@@
trains a diverse population of partners by a trajectory-based diversity measurement and a common best-response agent to these partners. Our implementation refers to the official demo [^3] https://colab.research.google.com/drive/1yIiw0Qs_d3O_JzGxtyvxB--jJ2M0h3Ki, where the diversity coefficient $\alpha$ is set to $4$.
@@
first trains a set of diverse partners and then trains a robust agent to coordinate well with these partners. The diversity is achieved by different random seeds and different checkpoints (i.e., different training stages). We maintain $5$ self-play partners initialized with $5$ different seeds (i.e., 1000–5000), respectively. Besides, we save the first checkpoint (i.e., a randomly initialized partner), the middle checkpoint (i.e., a half-trained partner), and the last checkpoint (i.e., a fully trained partner) to obtain the diverse population of partners.
@@
iis similar to FCP, which also first trains a set of diverse partners and then trains a robust agent to coordinate well with these partners. The diversity is achieved by an effective entropy-based loss term. We use the reference implementation of the authors for MEP [^4] https://github.com/ruizhaogit/maximum_entropy_population_based_training and employ the default parameter settings (i.e., the weight of the population entropy reward is $0.04$ for the FC layout and $0.01$ for other layouts).
COLE
constructs open-ended objectives from the perspective of graph theory to efficiently evaluate and identify cooperative incompatibility. We use the reference implementation of the authors for COLE [^5] https://github.com/liyang619/COLE-Platform/tree/COLE_training and employ the default parameter settings and population size $5$.
@@ V-C RQ1: Is the heterogeneity worth considering?
We first investigate if considering heterogeneity is valuable during the training process. To purely and clearly examine the importance of considering heterogeneity, we implement the simplest variant of MAZE, i.e., V-MAZE, which uses the fixed pairing strategy, updates the policies without considering the diversity term in the objective function, and does not use the archive for partner selection. That is, the only difference between V-MAZE and SP is that V-MAZE maintains two separate populations of agents and partners. We compare V-MAZE with SP and PP, which use self-play and single-population play, respectively. Note that we did not include the results of the methods (TrajeDi, FCP, and MEP), because they have other complex mechanisms, such as diversity-based updates to improve ZSC ability, and thus the performance difference cannot be purely attributed to the mechanism of considering heterogeneity.
@@
Table II: The reward (mean±$\pm$±std.) achieved by the compared algorithms when testing with different partners on different layouts. For each combination of layout and partner, the largest reward is bolded.
Layout Partner SP PP V-MAZE
H-CR Random 3.8 $\pm$ 2.1 5.0 $\pm$ 1.5 4.0 $\pm$ 0.7
Self-Play 40.4 $\pm$ 9.5 36.8 $\pm$ 6.4 50.3 $\pm$ 7.1
MAZE 100.0 $\pm$ 14.4 143.7 $\pm$ 17.1 124.5 $\pm$ 14.2
Human Proxy 61.4 $\pm$ 10.3 73.6 $\pm$ 10.3 69.5 $\pm$ 10.1
AA Random 30.6 $\pm$ 7.6 36.0 $\pm$ 9.1 48.0 $\pm$ 8.7
Self-Play 172.6 $\pm$ 10.4 204.8 $\pm$ 10.8 207.0 $\pm$ 20.9
MAZE 213.8 $\pm$ 17.0 223.0 $\pm$ 10.1 310.5 $\pm$ 52.6
Human Proxy 38.5 $\pm$ 10.0 43.9 $\pm$ 2.9 43.5 $\pm$ 6.0
AA-2 Random 28.0 $\pm$ 6.7 29.8 $\pm$ 7.0 52.0 $\pm$ 10.6
Self-Play 52.4 $\pm$ 10.4 79.6 $\pm$ 19.4 123.4 $\pm$ 29.9
MAZE 132.6 $\pm$ 29.7 152.8 $\pm$ 24.6 308.2 $\pm$ 34.3
Human Proxy 39.4 $\pm$ 6.2 33.8 $\pm$ 9.3 60.4 $\pm$ 14.3
FC Random 3.3 $\pm$ 2.2 3.0 $\pm$ 0.6 4.0 $\pm$ 1.9
Self-Play 74.6 $\pm$ 17.6 76.8 $\pm$ 25.2 99.0 $\pm$ 37.9
MAZE 96.5 $\pm$ 16.0 92.6 $\pm$ 11.6 108.5 $\pm$ 26.0
Human Proxy 21.3 $\pm$ 2.4 28.1 $\pm$ 13.3 18.3 $\pm$ 5.3
Grid-T Self-Play 79.3 $\pm$ 12.4 168.9 $\pm$ 17.2 242.4 $\pm$ 25.1
MAZE 197.3 $\pm$ 10.9 215.2 $\pm$ 20.8 358.6 $\pm$ 19.2
Grid-L Self-Play 74.2 $\pm$ 11.6 79.5 $\pm$ 14.3 201.0 $\pm$ 19.7
MAZE 184.8 $\pm$ 7.4 193.2 $\pm$ 8.9 304.5 $\pm$ 17.4
Average Rank 2.75 1.95 1.3
Table III: The reward (mean±$\pm$±std.) achieved by the compared algorithms when testing with different partners on different layouts. For each combination of layout and partner, the largest reward is bolded. The symbols ‘$+$’, ‘--’ and $\approx$’ indicate that the result is significantly superior to, inferior to, and almost equivalent to MAZE, respectively, according to the Wilcoxon rank-sum test with significance level 0.05.
Layout Partner PP TrajeDi FCP MEP COLE MAZE
H-CR Random 5.0 $\pm$ 1.5 -- 5.5 $\pm$ 0.7 -- 6.4 $\pm$ 1.1 $\approx$ 6.5 $\pm$ 0.7 $\approx$ 6.7 $\pm$ 0.6 $\approx$ 7.0 $\pm$ 0.7
Self-Play 36.8 $\pm$ 6.4 -- 87.6 $\pm$ 9.0 $\approx$ 121.6 $\pm$ 12.9 $+$ 85.8 $\pm$ 7.5 $\approx$ 88.9 $\pm$ 6.8 $\approx$ 97.5 $\pm$ 14.6
MAZE 143.7 $\pm$ 17.1 -- 153.8 $\pm$ 16.5 -- 140.4 $\pm$ 15.5 -- 129.8 $\pm$ 14.9 -- 138.2 $\pm$ 13.6 -- 183.7 $\pm$ 11.9
Human Proxy 73.6 $\pm$ 10.3 -- 81.0 $\pm$ 13.3 -- 101.2 $\pm$ 15.5 $\approx$ 90.2 $\pm$ 10.0 -- 102.4 $\pm$ 9.3 $\approx$ 118.6 $\pm$ 13.2
AA Random 36.0 $\pm$ 9.1 -- 49.3 $\pm$ 11.9 -- 50.4 $\pm$ 7.4 -- 63.5 $\pm$ 12.5 $\approx$ 59.7 $\pm$ 9.3 -- 74.4 $\pm$ 15.5
Self-Play 204.8 $\pm$ 10.8 $+$ 180.5 $\pm$ 15.5 $+$ 171.0 $\pm$ 11.5 $+$ 190.3 $\pm$ 10.1 $+$ 197.4 $\pm$ 12.6 $+$ 142.6 $\pm$ 20.3
MAZE 223.0 $\pm$ 10.1 -- 247.0 $\pm$ 27.1 -- 236.0 $\pm$ 12.9 -- 120.0 $\pm$ 12.4 -- 278.2 $\pm$ 11.4 -- 315.5 $\pm$ 10.1
Human Proxy 43.9 $\pm$ 2.9 -- 50.2 $\pm$ 6.0 -- 38.4 $\pm$ 4.2 -- 92.0 $\pm$ 29.5 $\approx$ 103.5 $\pm$ 14.1 $\approx$ 112.0 $\pm$ 19.1
AA-2 Random 29.8 $\pm$ 7.0 -- 45.5 $\pm$ 10.1 $\approx$ 43.5 $\pm$ 10.5 $\approx$ 58.4 $\pm$ 14.8 $\approx$ 46.3 $\pm$ 12.1 $\approx$ 56.5 $\pm$ 9.3
Self-Play 79.6 $\pm$ 19.4 -- 100.8 $\pm$ 12.5 -- 121.2 $\pm$ 15.5 $\approx$ 108.0 $\pm$ 29.9 $\approx$ 125.7 $\pm$ 14.7 $\approx$ 130.6 $\pm$ 18.5
MAZE 152.8 $\pm$ 24.6 -- 170.5 $\pm$ 11.1 -- 156.0 $\pm$ 15.3 -- 124.6 $\pm$ 29.3 -- 298.5 $\pm$ 14.8 -- 381.4 $\pm$ 13.7
Human Proxy 33.8 $\pm$ 9.3 -- 48.0 $\pm$ 6.9 -- 55.8 $\pm$ 15.2 -- 79.0 $\pm$ 11.2 -- 92.3 $\pm$ 14.1 -- 111.5 $\pm$ 13.7
FC Random 3.0 $\pm$ 0.6 -- 6.0 $\pm$ 0.7 -- 5.1 $\pm$ 0.7 -- 6.0 $\pm$ 1.0 -- 7.6 $\pm$ 2.0 $\approx$ 7.3 $\pm$ 0.4
Self-Play 76.8 $\pm$ 25.2 -- 80.8 $\pm$ 22.4 -- 79.5 $\pm$ 22.4 -- 82.3 $\pm$ 22.6 -- 125.6 $\pm$ 19.4 -- 147.0 $\pm$ 27.1
MAZE 92.6 $\pm$ 11.6 -- 115.6 $\pm$ 25.7 -- 119.8 $\pm$ 10.9 -- 95.2 $\pm$ 15.7 -- 131.2 $\pm$ 16.4 -- 164.0 $\pm$ 27.0
Human Proxy 28.1 $\pm$ 13.3 $\approx$ 25.4 $\pm$ 6.8 $\approx$ 26.9 $\pm$ 7.4 $\approx$ 28.6 $\pm$ 5.3 $\approx$ 35.7 $\pm$ 4.2 $+$ 25.0 $\pm$ 5.1
+//+/-$\approx$+ / - / ≈ 1/14/1 1/11/3 2/10/4 1/8/7 2/7/7
Average Rank 5.13 4.06 4.00 3.75 2.25 1.81
@@
The training curves are shown in Figure 6, reflecting the change in the average reward of the agents during the training phase. In all the six heterogeneous environments, V-MAZE achieves better performance clearly, showing the necessity of considering the heterogeneity and distinguishing the two players explicitly. Besides, we use a larger scale network to train SP and PP agents on the heterogeneous layout AA in Section V-G. With more generations (i.e., almost three times of V-MAZE), SP and PP still perform worse than V-MAZE. This indicates that the heterogeneous skills of both players cannot be mastered well at the same time even with enhanced representation ability and more data, further strengthening our conclusions. The test performance with unknown partners (defined in Section V-D in Table II align with the training curves, i.e., V-MAZE $>$ PP $>$ SP.